OpenStudy (anonymous):

What is a pair of conjugates with a product of 5

OpenStudy (anonymous):

(3-2)(3+2)

OpenStudy (anonymous):

A pair of conjugates is just (a+b)(a-b). A in this case being 3 and b being 2

OpenStudy (anonymous):

\[(2+i)(2-i)=2^2+1^2=5\] comes to mind

OpenStudy (anonymous):

If you want a complex conjugate.

OpenStudy (anonymous):

is that a conjugate

OpenStudy (anonymous):

@alrightmaths the conjugate of 1 is not 5

OpenStudy (anonymous):

3+2 and 3-2 are a pair of conjugates

OpenStudy (anonymous):

a+b and a-b

OpenStudy (anonymous):

i respectfully disagree

OpenStudy (anonymous):

the conjugate of \(a+bi\) is \(a-bi\) and the conjugate of \(a+\sqrt{b}\) is \(a-\sqrt{b}\) but the conjugate of a rational number is itself

OpenStudy (anonymous):

what would be the conjugate of 10? if you write \(10=8+2\) then would you say \(8-2\)? but what about \(7+3\)?

OpenStudy (anonymous):

I wouldn't say the conjugate of 10 was 8-2, I'd say the conjugate of 8+2 was 8-2

OpenStudy (anonymous):

Why does it make a difference or not whether the a and b are rational or not?

OpenStudy (anonymous):

would you agree that \(8+2=10\) ?

OpenStudy (anonymous):

No, pretty sure that's 14

OpenStudy (anonymous):

(\[(\sqrt{7}-\sqrt{2})(\sqrt{7}+\sqrt{2})\]